# FedDyMem: Efficient Federated Learning with Dynamic Memory and Memory-Reduce for Unsupervised Image Anomaly Detection

### v1

@@ VI-D1 Impact of Local Training Epochs
VI-D2 Effect of Projection Layer
The ablation study is presented in Table III shows the significant impact of incorporating a projection layer on the performance of the proposed FedDyMem framework. The baseline configuration used for comparison consists of memory-reduction and memory feature average aggregation mechanisms, without additional architectural enhancements. Specifically, introducing the projection layer improves both I-AUROC and P-AUROC metrics across all datasets. For example, the I-AUROC for Wafer increases from 67.69% to 68.93%, and the P-AUROC improves from 70.23% to 74.29%. These results demonstrate that the projection layer effectively enhances feature representations.
VI-D3 Effect of Memory Generator
The integration of the memory generator significantly improves the anomaly detection performance of FedDyMem across different datasets. For example, the memory generator increases the I-AUROC from 68.93% to 70.96% on the wafer dataset and from 74.10% to 81.26% on the PCB dataset. A similar trend is observed for P-AUROC, with increases from 74.29% to 77.01% on Wafer and from 89.26% to 94.97% on PCB. These improvements demonstrate the ability of the memory generator to effectively refine feature representations.
VI-D4 Effect of $K$-means
The addition of $K$-means to the FedDyMem framework resulted in a consistent improvement across all evaluated metrics. When combined with the projection layer and memory generator, $K$-means further improves I-AUROC, achieving increases such as from 70.96% to 71.96% for Wafer and from 81.26% to 83.34% for PCB. Similarly, P-AUROC improves significantly with values such as Wafer from 77.01% to 78.40% and PCB from 94.97% to 96.03%.
VI-E Qualitative Results
@@
Fig. 9 provides a qualitative comparison of the prediction results between FedAvg and the proposed FedDyMem method under an identical experimental setting. The FedDyMem approach exhibits a notable improvement in prediction quality, with substantially fewer noisy regions evident in the results. Following the methodology in [58], the anomaly score heatmaps are upsampled to match the spatial resolution of the input image and subsequently refined using Gaussian filtering with $\sigma=4$, ensuring smoother and more coherent boundaries. Furthermore, min-max normalization is applied to standardize the range of anomaly scores, facilitating clearer visualization and interpretation of the results.
VII Conclusion
@@
In this article, we propose a efficient federated learning with dynamic memory and memory-reduce, called FedDyMem to address the federated UAD problem. To tackle the challenges posed by multi-client data distribution bias in real-world industrial and medical scenarios, FedDyMem introduces a memory generator and a loss function based on distance metrics to dynamically generate high-quality memory banks. To enhance communication efficiency, a memory-reduce method is incorporated to decrease the size of the memory banks, thereby minimizing communication costs. During the aggregation phase, K-means is employed on the server to mitigate ambiguity and confusion across memory banks from different clients. Extensive experimental results demonstrate that the proposed method significantly outperforms existing baselines. Future work will focus on optimizing the model heterogeneity adaptation to further enhance the scalability and generalization of the proposed framework.

### v2

@@ 6.4 Ablation Study
Figure 11: Qualitative reconstruction results before and after applying the proposed memory-reduce operation on MVTec, PCB, and Brain.
6.5 Qualitative Results
@@
Fig. 10 provides a qualitative comparison of the prediction results between FedAvg and the proposed FedDyMem method under an identical experimental setting. The FedDyMem approach exhibits a notable improvement in prediction quality, with substantially fewer noisy regions evident in the results. Following the methodology in ([lee2022cfa]), the anomaly score heatmaps are upsampled to match the spatial resolution of the input image and subsequently refined using Gaussian filtering with $\sigma=4$, ensuring smoother and more coherent boundaries. Furthermore, min-max normalization is applied to standardize the range of anomaly scores, facilitating clearer visualization and interpretation of the results.
Table 5: Quantitative reconstruction attack results before and after memory-reduce.
Attack Dataset +Memory- SSIM PSNR FID
Methods Reduce
Deconvolution Inverting Method MVTec - 0.6350 25.9003 246.9916
0.2088 17.4787 415.9250
PCB - 0.5213 25.8625 281.5855
0.3379 12.1827 358.7053
Brain - 0.2273 11.0985 300.5543
0.0872 2.8752 338.1136
Collaborative inference attack MVTec - 0.0928 15.5441 300.9797
0.0179 11.4840 350.8446
PCB - 0.4860 13.9388 313.1274
0.0178 10.7200 418.3834
Brain - 0.1218 11.2046 365.9037
0.0038 6.5987 396.2980
Feature Reconstruction Attack MVTec - 0.2094 16.2157 231.2310
0.0869 15.3695 326.0596
PCB - 0.5391 17.4453 327.5368
0.2073 14.2317 370.4590
Brain - 0.5629 20.2312 240.1137
0.1558 7.8485 352.5879
6.6 Privacy Verification
To evaluate the privacy-preserving capability of FedDyMem, we conducted reconstruction attack experiments under representative adversarial settings. In these experiments, we considered the threat scenario in which an attacker attempts to recover original input images from the memory features exchanged between clients and the server. We applied three widely used attack methods, including the deconvolution-based inverting method ([dosovitskiy2016inverting]), the collaborative inference attack method ([he2019model]), and the stealthy feature-oriented reconstruction attack ([xu2024stealthy]). We evaluated the quality of the reconstructed images using three established metrics. The Structural Similarity Index (SSIM) and Peak Signal-to-Noise Ratio (PSNR) assess the perceptual and signal-level fidelity between the reconstructed and original images. The Fréchet Inception Distance (FID) measures the semantic distance in a high-level feature space and reflects how closely the generated samples resemble raw data. A higher SSIM and PSNR indicate stronger reconstruction fidelity and thus greater privacy leakage, while a lower FID suggests a closer match in feature distribution to the original images.
Table 5 presents the performance of each attack method across three datasets before and after applying the memory-reduce. Across all attack settings, memory-reduce consistently degrades both the reconstruction quality and the success rate of the adversaries. Specifically, we observe lower SSIM and PSNR values and higher FID scores after memory-reduce is applied, indicating reduced visual similarity and semantic fidelity between reconstructed and original images. Moreover, the ASR drops to 0% in all but one case, where it is reduced from 28.57% to 5.23%, highlighting the strong effectiveness of memory-reduce in mitigating reconstruction attacks. In addition to the quantitative results, we present qualitative comparisons in Fig. 11, which visualizes representative reconstructed samples under different settings. The reconstructions before applying memory-reduce retain identifiable structure or texture information from the original images, especially in regions with distinct contours. In contrast, after memory-reduce is applied, the reconstructed images exhibit significant degradation in visual fidelity. The overall shape, content, and fine-grained patterns become vague or distorted, demonstrating the effectiveness of memory-reduce in suppressing reconstruction attacks.
7 Conclusion
@@
In this article, we propose an efficient federated learning with dynamic memory and memory-reduce, called FedDyMem to address the federated UAD problem. To tackle the challenges posed by multi-client data distribution bias in real-world industrial and medical scenarios, FedDyMem introduces a memory generator and a loss function based on distance metrics to dynamically generate high-quality memory banks. To enhance both communication efficiency and privacy protection, a memory-reduce method is incorporated to compress memory features and reduce their mutual information with raw data. During the aggregation phase, K-means is employed on the server to mitigate ambiguity and confusion across memory banks from different clients. Extensive experimental results demonstrate that the proposed method significantly outperforms existing baselines. Future work will focus on optimizing the model heterogeneity adaptation to further enhance the scalability and generalization of the proposed framework.